OpenStudy (anonymous):

What is the inverse of y = 5 - 3x? Show all your steps

OpenStudy (anonymous):

write the same equation in term of y that is the inverse

Parth (parthkohli):

In inverse functions, the \(x\) replaces \(y\). Here,\[x = 5 - 3y\]Now solve for \(y\).

OpenStudy (anonymous):

Well, I kinda get the idea how to do this. Just want step by step. I have done similar problems.

OpenStudy (anonymous):

switch the x & y x = 5 - 3y subtract 5 from each side x - 5 = -3y divide by -3 on each side so (x-5)/3

Parth (parthkohli):

\[\begin{aligned}x-5&=-3y \\ {x -5 \over -3}&=y \end{aligned}\]

Parth (parthkohli):

Or,\[{-x + 5 \over 3} = y\]Or\[{5 - x \over 3} = y\]

OpenStudy (anonymous):

Ahh, so I am just switching x and y around?

Parth (parthkohli):

Yeah

OpenStudy (anonymous):

Gotcha, thanks ya'll! (: